OpenStudy (love_333):

Medal + fan! What is the volume of a square pyramid with base edges of 40cm and a slant of height of 29cm? A 11,200cm^3 B 15,467cm^3 C 16,800cm^3 D 33,600cm^3

OpenStudy (love_333):

I want the answer and how to work it out please, I haven't worked a problem like this before.

OpenStudy (love_333):

@Erak Ideas?

OpenStudy (erak):

So basically for a pyramid you need V = lwh/3 right? You are given l and w, because a square base with 40cm edges mean l and w = 40. For the height, you need the center line from the tip of the pyramid to the center of the base. Now try to imagine a right angle triangle, where your slant is the hypotenuse. Just ask and I will draw a bad drawing of this. So you would get your height by using Pythagoreas C^2 = A^2 + B^2 29^2 = 20^2 + B^2 <---that side is 20 because it's half of the base 841 = 400 + B^2 B^2 = 841 - 400 B^2 = 441 B = 21 <---your height Now just do V = lwh/3

OpenStudy (love_333):

AH, I see! Its A correct?

OpenStudy (erak):

yeah